A wheel 2 m in diameter rotates at 60 rpm. What is the acceleration of a point on the edge of the wheel? 7.
Answer #1

acceleration = v^2/r

v = w*r

a = (w*r)^2/r = w^2*r

w = 60 rpm = 6.28 rad/sec

r = 2 m

a = 6.28^2*2 = 78.88 m/sec^2